OpenStudy (anonymous):

Write an inequality for the sentence: twice a number, decreased by the quotient of that number and 2, is at least 12. plzz help

OpenStudy (anonymous):

if the number is say \(x\) then twice the number is \(2x\)

OpenStudy (anonymous):

the quotient of that number and two is \(\frac{x}{2}\) and if you decrease \(2x\) by \(\frac{x}{2}\) you get \[2x-\frac{x}{2}\]

OpenStudy (anonymous):

to say it is "at least 12" is tricky a bit do you think it means "greater than or equal to" or "less than or equal to"?

OpenStudy (anonymous):

greater than?

OpenStudy (anonymous):

yes exactly well "greater than or equal to "

OpenStudy (anonymous):

as in "you must be at least 12 to drink beer"

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

yw